These worded problems are killing me. Please assist me in making this a quadratic equation. Thank you.
It states:
Find the radius of a 12 cm high cylinder that has a surface area of 846cm^2.
Thank you so very much.

Answer by stanbon(75887)

Find the radius of a 12 cm high cylinder that has a surface area of 846cm^2.
-----------
Surface Area = 2(pi)rh
--------------------
846 = 2(pi)r*12
70.5 = 2(pi)r
r = 11.22 cm
================
Cheers,
Stan H.